Ben is 5 more than 3 times as old as henry. in 5 years the sum of thier ages will be 79. how old is henry now
Answer by Maths68(1474)

Let
Present age of Ben = b
Present age of Henry = h

Ben is 5 more than 3 times as old as Henry.
b=3h+5.............(1)
In 5 years
Ben will be = b+5
Henry will be = h+5
Then
The sum of their ages will be 79
(b+5)+(h+5)=79
b+5+h+5=79
b+h+10=79
b+h=79-10
b+h=69..............(2)
Put the value of b from (1) to (2)
3h+5+h=69
4h=69-5
4h=64
4h/4=64/4
h=16
Put the value of h in (1)
b=3h+5
b=3(16)+5
b=48+5
b=53


Present age of Ben = b = 53 years old
Present age of Henry = h = 16 years old